# Minimal Model for Reservoir Computing

### v2

@@ 6 Conclusion
We have demonstrated that a reservoir computing exists that emulates given coupled maps by constructing a modularised network. We have proposed a possible mechanism for the collapses of the emulation in reservoir computing. Such transitory behaviour is caused by either (i) a finite-time stagnation near an unstable chaotic set or (ii) a critical transition by the effective parameter drift. The essential problem in reservoir computing is determining why a finite-size reservoir computing with a randomly selected network can efficiently emulate dynamical systems with large degrees of freedom. Our approaches provides a minimal model for understanding reservoir computing, thereby providing better insights into the design of reservoir computing for practical applications. Problems of quenched random dynamical systems analysis for bifurcations, generalised synchronisation, and collapse of the emulation will be studied elsewhere.
Acknowledgements We acknowledge H. Suetani for useful discussion and comments. The research leading to these results has been partially funded by the Grant in Aid for Scientific Research (B) No. 21H01002, JSPS, Japan, and the London Mathematical Laboratory External Fellowship, United Kingdom.
@@ Appendix A Echo state property
The echo state property is a characteristic of a given reservoir dynamics with a network matrix $A$, where the reservoir dynamics converge to the same reservoir behavior as time tends to infinity, regardless of the choice of initial reservoir states Yildiz et al. (2012). As an empirical measure of the echo state property, the condition of the spectral radius of the network matrix $\rho(A)<1$ has been discussed. However, in our construction, we can set arbitrary spectral radii for perfect emulations of any coupled maps. In general, if multiple attractors exist, the echo state property does not hold Ceni et al. (2020).
@@ Appendix B Delay coordinate embedding by reservoir computing
Assuming that the training data is provided as one-dimensional time series $\{u(t)\}$ and the attractor of the dynamics generating the training data $\bm{y}(t+1)=\bm{g}(\bm{y}(t))$ is a compact manifold of dimension $d$, the non-autonomous reservoir computing given by Eq. (1) with $12d+$ 1 output neurons, and with a smooth activation function is embedding by the embedding theorem Takens (2006); Sauer et al. (1991). Both embedding and the reservoir computing unfold the training data into a high-dimensional space (or a reservoir), and take a projection of the high-dimensional structure. To see that with a simple example, we construct a $K=3$ reservoir computing $\cal{Q}$ with $1$-dimensional input neurons, $3$-dimensional output neurons, that embeds $1$-dimensional time series to $3$-dimensional state space. We assume that the training data are provided as $\{u(t)\}$ that is generated by a dynamical system whose effective dimension of stationary attractor is less than $3$. When the training data $\{u(t)\}$ is scaled to the dynamics very close to the origin, the activation function $\sigma(x)=\tanh(x)$ is an almost identical function as $\tanh(x)\simeq x~{}(x\simeq 0)$. We set $\gamma=1$, and
@@
$\left\{\begin{array}[]{l}x(t)=u(t)\\ \bm{y}(t)=W_{\rm{out}}^{*}\bm{r}(t)=\bm{r}(t)\\ r^{(1)}(t+1)=\tanh(x(t))\simeq x(t)\\ r^{(2)}(t+1)=\tanh(r^{(1)}(t))\simeq r^{(1)}(t)\\ r^{(3)}(t+1)=\tanh(r^{(2)}(t))\simeq r^{(2)}(t)\\ \end{array}\right..$ (30)
@@
$\left\{\begin{array}[]{l}y^{(1)}(t+3)=r^{(1)}(t+3)\simeq x(t+2)=u(t+2)\\ y^{(2)}(t+3)=r^{(2)}(t+3)\simeq x(t+1)=u(t+1)\\ y^{(3)}(t+3)=r^{(3)}(t+3)\simeq x(t)=u(t)\\ \end{array}\right.$ (31)
via a reservoir $\cal{Q}$. See Hart et al. (2020) for the general embedding theorem for non-autonomous echo state networks.

### v3

@@ 7 Conclusion
We have demonstrated that a reservoir computer exists that emulates given coupled maps by constructing a modularised network. We have proposed a possible mechanism for the collapses of the emulation in reservoir computing by introducing a measure of finite scale deviation. Such transitory behaviour is caused by either (i) a finite-time stagnation near an unstable chaotic set or (ii) a critical transition by the effective parameter drift. The essential problem in reservoir computing is determining why a finite-size reservoir computer with a randomly selected network can efficiently emulate dynamical systems with large degrees of freedom. Our approaches provides a minimal model for understanding reservoir computing, thereby providing better insights into the design of reservoir computer for practical applications. Problems in the quenched random dynamical systems, such as bifurcations, generalised synchronisation, and various types of the emulation collapse, will be studied elsewhere.
Acknowledgements
We acknowledge H. Suetani (Oita University) for useful discussion and comments. The research leading to these results has been partially funded by the Grant in Aid for Scientific Research (B) No. 21H01002, JSPS, Japan, and the London Mathematical Laboratory External Fellowship, United Kingdom.
@@ Appendix A Echo state property
The echo state property is a characteristic of a given reservoir dynamics with a network matrix $A$, where the reservoir dynamics converge to the same reservoir behavior as time tends to infinity, regardless of the choice of initial reservoir states [24]. As an empirical measure of the echo state property, the condition of the spectral radius of the network matrix $\rho(A)<1$ has been discussed. However, in our construction, we can set arbitrary spectral radii for perfect emulations of any coupled maps. In general, if multiple attractors exist, the echo state property does not hold [25].
@@ Appendix B Delay coordinate embedding by reservoir computing
Assuming that the training data is provided as one-dimensional time series $\{u(t)\}$ and the attractor of the dynamics generating the training data $\bm{y}(t+1)=\bm{g}(\bm{y}(t))$ is a compact manifold of dimension $d$, the non-autonomous reservoir computer given by Eq. (1) with $12d+$ 1 output neurons, and with a smooth activation function is embedding by the embedding theorem [14, 15]. Both embedding and the reservoir computing unfold the training data into a high-dimensional space (or a reservoir), and take a projection of the high-dimensional structure. To see that with a simple example, we construct a $K=3$ reservoir computer $\cal{Q}$ with $1$-dimensional input neurons, $3$-dimensional output neurons, that embeds $1$-dimensional time series to $3$-dimensional state space. We assume that the training data are provided as $\{u(t)\}$ that is generated by a dynamical system whose effective dimension of stationary attractor is less than $3$. When the training data $\{u(t)\}$ is scaled to the dynamics very close to the origin, the activation function $\sigma(x)=\tanh(x)$ is an almost identical function as $\tanh(x)\approx x~{}(x\approx 0)$. We set $\gamma=1$, and
@@
$\left\{\begin{array}[]{l}x(t)=u(t)\\ \bm{y}(t)=W_{\rm{out}}^{*}\bm{r}(t)=\bm{r}(t)\\ r^{(1)}(t+1)=\tanh(x(t))\approx x(t)\\ r^{(2)}(t+1)=\tanh(r^{(1)}(t))\approx r^{(1)}(t)\\ r^{(3)}(t+1)=\tanh(r^{(2)}(t))\approx r^{(2)}(t)\\ \end{array}\right..$ (33)
@@
$\left\{\begin{array}[]{l}y^{(1)}(t+3)=r^{(1)}(t+3)\approx x(t+2)=u(t+2)\\ y^{(2)}(t+3)=r^{(2)}(t+3)\approx x(t+1)=u(t+1)\\ y^{(3)}(t+3)=r^{(3)}(t+3)\approx x(t)=u(t)\\ \end{array}\right.$ (34)
via a reservoir $\cal{Q}$. See [13] for the general embedding theorem for non-autonomous echo state networks.